OpenStudy (anonymous):

can someone please explain to me how to find the answer------ Simplify 1/4(g − 16)

OpenStudy (ash2326):

We have \[\frac 14 (g-16)\] just multiply both terms by 1/4 you'd get \[\frac g 4- \frac{16}{4}\] Could you simplify now?

OpenStudy (anonymous):

how would you multiply the terms by 1/4

OpenStudy (ash2326):

it's outside the bracket, when I open it, I've to multiply the terms inside by 1/4 \[A(B+C)=A\times B+A\times C\]

OpenStudy (anonymous):

i dont want the answer i want to know how to do it so i can give the answer

OpenStudy (anonymous):

yes i know but how would you multiply a term by the number 1/4??

OpenStudy (ash2326):

multiplying by 1/4 is the same as dividing it by 4 so divide both the terms by 4

OpenStudy (ash2326):

do you get my point?

OpenStudy (anonymous):

ooooo ok so when i simplify would it become 1/4g + 4?

OpenStudy (ash2326):

yeah:) you're right

OpenStudy (anonymous):

THANK u soo much

OpenStudy (ash2326):

glad to help=)